OpenStudy (anonymous):

Help checking work with De Moivre's theorem? Let z=13+7i and w=3[cos(1.43)+isin(1.43)] A) Convert z to polar form My work: r=√13^2+7^2=√218=14.765 Θ=tan^-1 7/13=0.494 (14.765, 0.494) B) Calculate zw using De Moivre's theorem My work: 3 cos 1.43=0.421 3 i sin 1.43=2.970 i Multiply both by (13+7i) 5.473+2.947i+-20.79+38.61i=-15.317+41.557i C) Calculate z/w using De Moivre's theorem My work: (13+7i)/(0.421)=30.879+16.627i (13+7i)/(2.970i)=2.357-4.377i 33.236+12.25i

OpenStudy (irishboy123):

it says "using de Moivre" so maybe you should multiply/divide r's and add/subtract \(\theta\)'s before putting into rectangular form

OpenStudy (anonymous):

How would you do that? I'm not sure how to utilize de Moivre's here...

OpenStudy (anonymous):

multiply \(3\times \sqrt{218}\) for the modulus (absolute value) then add the angles

OpenStudy (anonymous):

btw your answer is wrong there

OpenStudy (anonymous):

i mean the multiplication lets assume \[13+7i=\sqrt{218}\cos(.494)+i\sin(.494))\]

OpenStudy (anonymous):

then to get the product, you multiply the numbers out front and add the angles that is demoivre

OpenStudy (anonymous):

I know it's a lot to ask, but could you walk me through it? I'm not sure what to multiply and what to add.

OpenStudy (anonymous):

oh damn, your angle is wrong too

OpenStudy (anonymous):

lets check all

OpenStudy (anonymous):

\[13+7i\] in polar form you need \(r=\sqrt{13^2+7^2}=14.765\) rounded so that part is right

OpenStudy (anonymous):

\[\theta=\tan^{-1}(\frac{7}{13})=0.494\] ok that part is right too, sorry

OpenStudy (anonymous):

so \[13+7i=14.765\left(\cos(.494)+i\sin(.494)\right)\]

OpenStudy (anonymous):

then \[14.765\left(\cos(.494)+i\sin(.494)\right)\times 3\left(\cos(1.43)+i\sin(1.43)\right)\] is the last job

OpenStudy (anonymous):

it will be \[14.765\times 3\left(\cos(.494+1.43)+i\sin(.494+1.43)\right)\]

OpenStudy (anonymous):

multiply the modulus, add the angles

OpenStudy (anonymous):

you still have to do the arithmetic ( use a calculator) but that is what you get before doing it

OpenStudy (anonymous):

let me know if it is not clear

OpenStudy (anonymous):

Wait, so 14.765x3(cos(.494+1.43)+i sin (.494+1.43)) is z*w, right?

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

still have to do the arithmetic though

OpenStudy (anonymous):

What would the equation to find z/w be? Is it 14.765x3(cos(.494-1.43)+i sin (.494-1.43))?

OpenStudy (anonymous):

close

OpenStudy (anonymous):

subtract the angles is right but divide the modulus

OpenStudy (anonymous):

So 14.765/3(cos(.494-1.43)+i sin (.494-1.43))? Sorry for all the questions!

OpenStudy (anonymous):

no problem and yes

OpenStudy (anonymous):

don't forget to do the arithmetic, or your teacher will think you are real lazy

OpenStudy (anonymous):

I got 2.9186-3.9629i for z/w and -15.321+41.561i for z*w.

OpenStudy (anonymous):

i will take your word for it

OpenStudy (anonymous):

Alright, sweet, thank you so much! I've been stuck on this problem for a while.

OpenStudy (anonymous):

it didn't say to rewrite in standard form i would have left it in polar form

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Wait is leaving it in polar form just 44.295(cos(1.924)+i sin (1.924)) ?